OpenStudy (anonymous):

Find the depressed polynomial of (x^3+4x^2-3x-5)÷(x+3)

OpenStudy (campbell_st):

Q(x) = x^2 +x -6 R(x) = 13/(x+3) P(x) = (x+3)(x^+x-6) + 13/(x+3)